OpenStudy (anonymous):

Help, medal Model the function rule y = -2x - 3 with a table of values.

OpenStudy (anonymous):

@CrazySouthernGurl14 @jigglypuff314

jigglypuff314 (jigglypuff314):

erm, I'm not entirely sure how to do this but I'm guessing you would make a table x y 1 2 3 ect then to get y values you would plug in the corresponding value of x to solve for y for example when x = 1 y = -2(1) - 3 = -2 - 3 = -5 therefore y = -5 when x = 1

OpenStudy (anonymous):

arrgg okay thank you, u think you could help me these other 3 questions?

jigglypuff314 (jigglypuff314):

I can try :)

OpenStudy (anonymous):

ok :) ty

jigglypuff314 (jigglypuff314):

m = 20 times 5 divided by 4

OpenStudy (anonymous):

25?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs.

jigglypuff314 (jigglypuff314):

I'm getting lazy sorry :P http://answers.yahoo.com/question/index?qid=20111205192422AAyXMmy

OpenStudy (anonymous):

lol okay ty :)

jigglypuff314 (jigglypuff314):

you're welcome :)